Here is my question: A banker invested a portion of $10,000 at 5% interest per year and the remainder at 4% interest per year. If the banker received $470 in interest after 1 year, how much was invested at each rate?
I really don't get this, PLEASE HELP!!!!

Answer by elima(1433)

Let x be the amount he invested at 5%,
10,000-x=amount invested at 4%
so we have;
5%x+(10,000-x)4%=470
5%x+400-4%x=470
400+1%x=470
1%x=70
x=7000
So if $7,000 was invested at 5%;
then $7,000-10,000=$3,000
$3,000 wasinvested at 4%
Hope you understand
=)